OpenStudy (anonymous):

The Greasy Grab Bag claims that customers can make 20 different meal combinations from the menu items. If the Greasy Grab Bag has 4 sandwiches, 3 sides, and 2 drinks, which of the following is correct? The Greasy Grab Bag is understating the number of meal combinations. The Greasy Grab Bag is stating the correct number of meal combinations. The Greasy Grab Bag is overstating the number of meal combinations.

OpenStudy (anonymous):

The Greasy Grab Bag is understating the number of meal combinations. sw1 + s1 + 1d sw1 + s1 + 2d sw1 + s2 + 1d sw1 + s2 + 2d = 6 possible sw1 + s3 + 1d sw1 + s3 + 2d sw2 + s1 + 1d sw2 + s1 + 2d sw2 + s2 + 1d sw2 + s2 + 2d = 6 possible sw2 + s3 + 1d sw2 + s3 + 2d sw3 + s1 + 1d sw3 + s1 + 2d sw3 + s2 + 1d sw3 + s2 + 2d = 6 possible sw3 + s3 + 1d sw3 + s3 + 2d sw4 + s1 + 1d sw4 + s1 + 2d sw4 + s2 + 1d sw4 + s2 + 2d = 6 possible sw4 + s3 + 1d sw4 + s3 + 2d 6 + 6 + 6 + 6= 24 24 is greater than 20